# An effective van der Corput-type method in higher dimensions

### v1

@@ 1. Introduction
(1.12) $I(\lambda)\approx\sum_{k}\sum_{l=1}^{l_{k}}a_{k,l}\lambda^{-p_{k}}\ln^{l-1}\lambda,$
@@ 5. Proof of the Main Theorem
(5.6) $f\circ\pi(\sigma)(y)=y_{1}^{l(\alpha^{1})}\cdots y_{n}^{l(\alpha^{n})}f_{\sigma}(y_{1},\cdots,y_{n}),$
and $f(0)\neq 0$.
The exceptional set $\pi_{\sigma}^{-1}(0)$ is a union of coordinate hyperplanes:
(5.7) $\pi_{\sigma}^{-1}(0)=\bigcup_{j\in I}\{y_{j}=0\}$
where $I\subset\{1,\cdots,n\}$ collects the indices corresponding to the compact face $\gamma$ associated with $\left\{\alpha^{1},\cdots,\alpha^{s}\right\}$. Without loss of generality, we may relabel coordinates so that $I=\{1,\cdots,s\}$, and denote
(5.8) $T_{I}=\left\{(y_{1},\cdots,y_{n}):y_{j}=0,j\in I\right\}.$
Consequently, we may decompose the pullback of $f$ as
(5.9) $f\circ\pi(\sigma)(y)=y_{1}^{l(\alpha^{1})}\cdots y_{s}^{l(\alpha^{s})}y_{s+1}^{l(\alpha^{s+1})}\cdots y_{n}^{l(\alpha^{n})}\left(\bar{f_{\sigma}}(y_{s+1},\cdots,y_{n})+y^{p}h(y)\right),$
where $\bar{f_{\sigma}}(0)\neq 0,p=(p_{1},\cdots,p_{n})$ satisfies $p_{j}>0$ for all $j$, and $y^{p}h(y)$ consists of terms of strictly higher order relative to the face $\gamma$. A key property of $\bar{f_{\sigma}}$ is the following. Let
(5.10) $w=(0,\cdots,0,w_{s+1},\cdots,w_{n})\in\mathbb{R}^{n}\backslash\pi(\sigma)^{-1}(0),$
if $\bar{f_{\sigma}}(w)=0$, then there exists $w_{j}\neq 0,j>s$ such that $\partial_{y_{j}}\bar{f_{\sigma}}(w)\neq 0$. This can be found in [Var76].
@@
(5.11) $I(\lambda)=\int_{V}e^{i\lambda y_{1}^{l(\alpha^{1})}\cdots y_{s}^{l(\alpha^{s})}y_{s+1}^{l(\alpha^{s+1})}\cdots y_{n}^{l(\alpha^{n})}\left(\bar{f_{\sigma}}(y_{s+1},\cdots,y_{n})+y^{p}h(y)\right)}\left|y_{1}\right|^{\left|\alpha^{1}\right|-1}\cdots\left|y_{n}\right|^{\left|\alpha^{n}\right|-1}\varphi_{k}(y)\varphi\circ\pi(\sigma)(y)\,\mathrm{d}y$
@@
(5.12) $1=\sum_{j}\psi_{j}+\sum_{l}\eta_{l}.$
Each $\psi_{j}$ is supported where $\bar{f_{\sigma}}(y_{s+1},\cdots,y_{n})$ does not change sign, while the union of the supports of the $\eta_{l}$ contains all zeroes of $\bar{f_{\sigma}}(y_{s+1},\cdots,y_{n})$. Inserting this decomposition into $I(\lambda)$ we obtain corresponding integrals $I_{j}(\lambda)$ and $I_{l}(\lambda)$. It suffices to estimate each piece separately.
On the support of $\psi_{k}$, the function $\bar{f_{\sigma}}(y)$ is bounded from zero. Thus $\bar{f_{\sigma}}$ and the higher order term may be absorbed into any one of the variables of $y_{s+1},\cdots,y_{n}$. For definiteness, we introduce
(5.13) $u_{1}=y_{1},\cdots,u_{s+1}=\epsilon y_{s+1}\left(\bar{f_{\sigma}}(y_{s+1},\cdots,y_{n})+y^{p}h(y)\right),\cdots,u_{n}=y_{n}$
where $\epsilon=\pm 1$, the choice is determined by the sign of $y_{s+1}^{l(\alpha^{s+1})}$. This is an invertible change of variables provided the support of $\psi_{j}$ is chosen sufficiently small. After this substitution, the phase becomes a pure monomial, then
@@
On the support of $\eta_{l}$, all zeroes lie inside, but by assumption the gradient of $\bar{f_{\sigma}}(y_{s+1},\cdots,y_{n})$ does not vanish in the support. Shrinking the support of $\eta_{l}$ if necessary, we may assume $\partial_{y_{s+1}}\bar{f_{\sigma}}\neq 0$ and $y_{s+1}\neq 0$ throughout this neighborhood. Therefore we can introduce the change of variables
(5.17) $u_{1}=y_{1},\cdots,u_{s+1}=\bar{f_{\sigma}}(y_{s+1},\cdots,y_{n})+y^{p}h(y),\cdots,u_{n}=x_{n}.$
@@
Fix $x_{1},\cdots,x_{s},x_{s+2},\cdots,x_{n}$, and perform $N$ integrations by parts with respect to $x_{s+1}$, choosing $N$ such that
@@
Summing over all pieces then completes the proof.

### v2

@@ 1. Introduction
(1.12) $I(\lambda)\sim\sum_{k}\sum_{l=1}^{l_{k}}a_{k,l}\lambda^{-p_{k}}\ln^{l-1}\lambda,$
@@ 5. Proof of the Main Theorem
(5.6) $f\circ\pi(\sigma)(y)=y_{1}^{l(\alpha^{1})}\cdots y_{n}^{l(\alpha^{n})}f_{\sigma}(y),$
and $f_{\sigma}(0)\neq 0$.
The exceptional set $\pi(\sigma)^{-1}(0)$ is a union of coordinate hyperplanes:
(5.7) $\pi(\sigma)^{-1}(0)=\bigcup_{j\in I}\{y_{j}=0\}$
where $I\subset\{1,\cdots,n\}$ collects the indices corresponding to a compact face $\gamma$. Without loss of generality, we may assume $I=\{1,\cdots,s\}$, so that $\gamma$ is associated with the set of skeleton vectors $\left\{\alpha^{1},\cdots,\alpha^{s}\right\}$.
For any set $J\subset\{1,\cdots,n\}$, define
(5.8) $T_{J}=\left\{(y_{1},\cdots,y_{n}):y_{j}=0,j\in J\right\}$
which satisfies
(5.9) $\pi(\sigma)(T_{J})=0.$
It is clear that $T_{J}\subset\pi(\sigma)^{-1}(0)$.
The $\mathbb{R}$-nondegeneracy of $f_{\sigma}$ yields a fundamental structural property of $f_{\sigma}(y)$: the set
(5.10) $\left\{y\in\pi^{-1}(0):f_{\sigma}(y)=0\right\}$
is nonsingular. Consequently, for every point $P\in T_{J}$, there exists an index $j^{\prime}\notin J$ such that
(5.11) $\frac{\partial f_{\sigma}}{\partial y_{j^{\prime}}}(P)\neq 0.$
This fact originates in Varchenko’s analysis [Var76], and a detailed exposition may be found in [Nos12].
@@
(5.12) $I(\lambda)=\int_{V}e^{i\lambda y_{1}^{l(\alpha^{1})}\cdots y_{n}^{l(\alpha^{n})}f_{\sigma}(y)}\left|y_{1}\right|^{\left|\alpha^{1}\right|-1}\cdots\left|y_{n}\right|^{\left|\alpha^{n}\right|-1}\varphi_{k}(y)\varphi\circ\pi(\sigma)(y)\,\mathrm{d}y$
@@
(5.13) $1=\sum_{k}\psi_{k}+\sum_{l}\eta_{l}.$
Each $\psi_{k}$ is supported where $f_{\sigma}(y)$ does not change sign, while the union of the supports of the $\eta_{l}$ contains all zeroes of $f_{\sigma}(y)$. Inserting this decomposition into $I(\lambda)$ we obtain corresponding integrals $I_{k}(\lambda)$ and $I_{l}(\lambda)$. It suffices to estimate each piece separately.
On the support of $\psi_{k}$, the function $f_{\sigma}(y)$ is bounded from zero. Thus $f_{\sigma}(y)$ may be absorbed into any one of the variables of $y_{s+1},\cdots,y_{n}$. For definiteness, we introduce
(5.14) $u_{j}=y_{j}\quad(j\neq s+1),\quad u_{s+1}=y_{s+1}f_{\sigma}(y).$
This is an invertible change of variables provided the support of $\psi_{k}$ is chosen sufficiently small. After this substitution, the phase becomes a pure monomial, then
@@
Since the region of integration is localized in a thin neighborhood of the union of coordinate hyperplanes $\pi(\sigma)^{-1}(0)$, we may assume that each $\eta_{l}$ is supported in a small neighborhood of a point
(5.18) $w^{l}=(w_{1}^{l},\cdots,w_{n}^{l})\in T_{J_{l}}\bigcap\{y:f_{\sigma}(y)=0\}.$
By the discussion above, we may, without loss of generality, take the index $j^{\prime}$ appearing earlier to be $s+1\notin J_{l}$, so that
(5.19) $\partial_{y_{s+1}}f_{\sigma}(w^{l})\neq 0\quad\text{and }\quad w_{j}^{l}\neq 0\text{ for }j\notin J_{l}.$
Thus, by shrinking the support of $\eta_{l}$ if necessary, we may assume that for all $y=(y_{1},\cdots,y_{n})\in\operatorname*{supp}(\eta_{l})$, one has $y_{s+1}\neq 0$. Consequently, we may introduce the following change of variables:
(5.20) $u_{j}=y_{j}\quad(j\neq s+1),\quad u_{s+1}=y_{s+1}^{l(\alpha^{s+1})}f_{\sigma}(y).$
@@
Fix $u_{1},\cdots,u_{s},u_{s+2},\cdots,u_{n}$, and perform $N$ integrations by parts with respect to $u_{s+1}$, choosing $N$ such that
@@
Summing over all pieces then completes the proof. ∎